ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60J80, 60G57, 60k35 Fleischmann, Klaus, WIAS, Berlin Super-Brownian motion with a super-Brownian catalyst A brief survey is given on results about the catalytic super-Brownian motion in R^{d}, d ≤ 3, where branching is controlled by the collision local time of an intrinsic Brownian particle with an ordinary super-Brownian motion. The talk is based on the papers listed below. D. A. Dawson and K. Fleischmann. A continuous super-Brownian motion in a super-Brownian medium. Journ. Theoret. Probab. , 10(1):213–276, 1997. D. A. Dawson and K. Fleischmann. Longtime behavior of a branching process controlled by branching catalysts. Stoch. Process. Appl. , 71(2):241– 257, 1997. J. -F. Delmas and K. Fleischmann. Collision local time of catalyst and reactant for a catalytic super-Brownian motion. In preparation, 1998. A. M. Etheridge and K. Fleischmann. Persistence of a two-dimensional super-Brownian motion in a catalytic medium. Probab. Theory Relat. Fields, 110(1):1–12, 1998. K. Fleischmann and A. Klenke. Smooth density field of catalytic super-Brownian motion. WIAS Berlin, Preprint No. 331; Ann. Appl. Probab. (submitted), 1997.